H=1/2g^2+f
Express f in terms of g and h

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

H=1/2g^2+f
Express f in terms of g and h
------
f = H - (1/2)g^2
----------------
Cheers,
Stan H.
------------

Answer by ikleyn(53763)

.
H = 1/2g^2+f
Express f in terms of g and h
~~~~~~~~~~~~~~~~~~~

Move the term 1/2h^2 to the left side with the opposite sign.